OpenStudy (anonymous):

How would I find the slope to (-7,-8) (-3,-1)

OpenStudy (anonymous):

rise/ run

OpenStudy (anonymous):

^^

OpenStudy (anonymous):

\[\frac{-1 -(-8)}{-3-(-7)}\]

OpenStudy (anonymous):

(y1-y2)/(x1-x2)=m(slope)

OpenStudy (anonymous):

@moneybird> I wrote that answer down and the answer came out negative for me but its really a positive...what am I doing wrong?

OpenStudy (anonymous):

when you minus a negative number, it means plus

OpenStudy (anonymous):

slope = -1+8 7 ----- =---- -3+7 4